OpenStudy (anonymous):

What is the simplified form of 2x-6/x-3? Im not quite sure where to start? Help please :)

OpenStudy (anonymous):

Would I start by factoring the top fraction with 2, as in 1(x-3)?

OpenStudy (anonymous):

\[\frac{ 2(x-3)}{ x-3 }\] = 2 x not equal to 3

OpenStudy (anonymous):

Does 2 not divide into 1?

OpenStudy (anonymous):

Hi sonny :) First try to factor out numerator 2x - 6 so we will get 2 ( x- 3) Then cancel the common term from numerator and denominator ...

OpenStudy (anonymous):

it's but 2/1 =2

OpenStudy (anonymous):

Alright, thank you both for your contributions. Always appreciated! So 2 is the answer and the restriction is 3 because x-3=0? :)

OpenStudy (anonymous):

Correct !! have a good day :)

OpenStudy (anonymous):

You as well c:

OpenStudy (anonymous):

yes, good luck